# Advances in Feed-Forward 3D Reconstruction and View Synthesis: A Survey

### v3

@@ 2.3.1 Gaussian Map
A Gaussian map refers to a 2D-based representation of 3D Gaussians, where each 2D location encodes a 3D Gaussian. As a pioneering effort, Splatter Image [74] employs a U-Net encoder-decoder architecture [78] to predict pixel-aligned 3D Gaussians for single-view 3D object reconstruction as illustrated in Fig. 5. To improve the reconstruction quality, subsequent methods leverage generic scene priors learned from large-scale datasets. GRM [79] directly maps input image pixels to pixel-aligned 3D Gaussians for feed-forward object reconstruction, based on LRM [10] that learns general reconstruction priors from large-scale datasets of 3D objects; Flash3D [80] further introduces a high-quality depth predictor as prior to achieve single-view scene-level reconstruction. Concurrently, GS-LRM [81] incorporates Transformer-based LRM to formulate per-pixel Gaussian prediction as a sequence-to-sequence mapping, achieving remarkable performance across both objects and scenes. This line of research has been further advanced by several works yuelei here: eFreeSplat [82] leverages a large vision transformer encoder [52] as 3D priors for Gaussian image prediction; Long-LRM [83] that combines Mamba2 blocks [84] with Transformer layers to handle long sequences of input images; and FreeSplatter [85] achieves pose-free Gaussian map prediction. However, these works are limited to reconstructing existing image observations without generative capabilities. To overcome this limitation, LGM [86] introduces pre-trained diffusion models [87, 12, 88, 89] to generate multiview images, which are then used by large multiview Gaussian models for multiview Gaussian prediction. Wonderland [90] further enhances this approach by leveraging a pre-trained video diffusion model [91] to generate informative video latents from a single image for 3DGS prediction. In addition to the methods above, another line of work focuses on the geometric quality of 3D scene reconstruction by incorporating geometric designs or priors, such as epipolar constraints, cost volumes, and pre-trained reconstruction models.
@@
Epipolar-based Methods. As a pioneering epipolar-based method, PixelSplat [5] leverages an epipolar line to resolve the scale ambiguity issue and capture cross-view features, estimating a probabilistic depth distribution as 3D Gaussian positions. Although PixelSplat is effective in regions strongly correlated with the input observations, it struggles in areas of high uncertainty, leading to blurry or failed reconstructions. To address this, LatentSplat [92] combines the strengths of regression-based and generative approaches to obtain high-quality reconstructions in uncertain areas. Building further on PixelSplat, GGRt [93] introduces a joint learning framework that integrates pose estimation with a 3DGS prediction network to enable pose-free 3D Gaussian prediction.
@@
Cost Volume-based Methods. A key limitation of PixelSplat is the inherent ambiguity and unreliability in mapping image features to depth distributions, resulting in suboptimal geometry reconstruction. To address this problem, MVSplat [94] and MVSGaussian [95] adopt a plane-sweeping-based cost volume to facilitate multiview Gaussian image prediction, leveraging cross-view feature similarities to improve depth estimation. However, these methods heavily depend on precise multiview feature matching, which becomes particularly challenging in scenes with occlusions, low texture, or repetitive patterns. To address this issue, TranSplat [96] introduces a depth-aware deformable matching transformer to generate a depth confidence map to facilitate multiview feature matching. Similarly, DepthSplat [97] uses a multiview depth model that leverages pre-trained monocular depth features to enhance the feed-forward 3DGS reconstruction. This line of research is further extended by several works. HiSplat [98] introduces hierarchical representations to capture large-scale structures and fine texture details; PanSplat [99] builds a hierarchical spherical cost volume for 4K panorama view synthesis; MVSplat360 [100] adapts MVSplat to support 360° novel view synthesis for large-scale scenes; and LongSplat [101] designs an online updating framework for long-sequence image streams.
@@
Pre-trained Reconstruction Models. With the advent of feed-forward pointmap reconstruction methods [1, 4, 62] as mentioned in Sec. 2.2, several feed-forward 3DGS methods directly use pre-trained models to generate dense pointmaps for 3D Gaussian reconstruction. For example, Splatt3R [76] builds on the large-scale pretrained foundation 3D MASt3R model [4] by integrating a Gaussian decoder, enabling pose-free feed-forward 3DGS. NoPoSplat [102] also uses MASt3R as the backbone and predicts 3D Gaussians in a canonical space without ground-truth camera poses and depth. SmileSplat [103] instead uses DUSt3R as the backbone to predict Gaussian surfels with a multi-head Gaussian regression decoder. SelfSplat [104] unifies Gaussian prediction with self-supervised learning of depth and camera poses, enabling simultaneous prediction of geometry, pose, and Gaussian attributes. However, methods relying on DUSt3R and MASt3R inherit the limitation of pairwise inputs, which restricts their scalability. Thus, PREF3R [105] adopts Spann3R [62] as the geometric prior, which introduces a spatial memory network to handle multiview images. Except for feed-forward pointmap prediction models, pre-trained 3D diffusion models [25] have also been explored as a geometric prior for 3D Gaussian prediction [106].
@@ 2.3.2 Gaussian Volume
Gaussian volume [75] represents 3D with Gaussian voxel grids, where each voxel comprises multiple Gaussian primitives. A typical feed-forward 3DGS method using a Gaussian volume representation is LaRa [75], which aims to reduce the heavy training cost associated with 360° bounded radiance field reconstruction. As shown in Fig. 5 (b), it first builds 3D features and embedding volumes and then leverages a volume transformer to reconstruct a Gaussian volume, enabling progressively and implicitly feature matching that leads to higher quality results and faster convergence. Building on the notion of structured volumetric Gaussians, GaussianCube [107] proposes a structured and explicit radiance representation for 3D object generation from a single image. In parallel, QuickSplat [108], based on G3R [109], predicts voxel-level features using neural networks, which are subsequently decoded into Gaussian primitives with MLPs for surface reconstruction. To further enrich volumetric representations, GD [110] builds upon LaRa and introduces a generative densification that exploits prior knowledge from large multiview datasets and densifies feature representations from feed-forward 3DGS. SCube [111] advances this direction toward large-scale scene reconstruction by proposing VoxSplat, a high-resolution sparse-voxel Gaussian representation generated via a hierarchical latent diffusion model conditioned on sparse posed images. Most recently, AnySplat [112] constructs volumetric Gaussians by voxelizing the Gaussian maps from variable numbers of input views, yielding a unified and efficient model capable of handling both sparse and dense views.
To mitigate the high computational cost of the 3D volume, an efficient triplane structure [42, 43] is also explored for Gaussian volume prediction. It is typically constructed by predicting a triplane representation first and then leveraging the latent triplane features to decode 3D Gaussians, as illustrated in Fig. 5. For example, Triplane-Gaussian [77] leverages several transformer-based networks pre-trained in large-scale datasets to build a Gaussian triplane, enabling high-quality single-view 3D reconstruction. AGG [113] also mixes triplane and 3D Gaussians, which first represents scene textures as triplane and then decodes 3D Gaussians from triplane-based texture features queried by 3D locations.
@@ 2.4.1 Mesh
Meshes are compatible with various graphics pipelines and have gained significant attention in feed-forward 3D reconstruction in recent years. For instance, Pixel2Mesh [114] generates a 3D mesh from a single image by extracting features with a 2D CNN and progressively deforming an initial mesh; Mesh R-CNN [115] introduces a voxel branch based on Mask R-CNN [116] to predict a coarse cubified mesh followed by further refinement.
This field has been further advanced by the development of diffusion models. For example, One-2-3-45 [117] builds on the diffusion-based model to generate multiview images, which are then processed by a generalizable surface reconstruction module [118] for mesh reconstruction; One-2-3-45++ [119] further utilizes a 3D diffusion-based module conditioned on multiple views to generate a textured mesh in a coarse-to-fine manner. To improve geometry consistency, Wonder3D [87] introduces a cross-domain diffusion model to generate multiview-consistent normal maps and RGB images simultaneously, which can be used for mesh reconstruction. Similarly, Unique3D [120] employs a multiview diffusion model alongside a normal diffusion model to generate multiview-consistent images and normal maps, which are fed into a fast and consistent mesh reconstruction module.
In addition, several methods are proposed to take advantage of large reconstruction model [10] to achieve high-quality mesh reconstruction. For example, MeshLRM [121] integrates differentiable surface extraction and rendering into a large reconstruction model, enabling direct 3D mesh generation; InstantMesh [122] utilizes a transformer-based large reconstruction model to generate a high-quality 3D mesh from multiview images; MeshFormer [11] leverages 3D voxel representations and combines transformers with 3D convolutions to improve 3D mesh geometry.
To generate artist-created meshes with high-quality topology, several methods draw inspiration from large language models, treating 3D meshes as sequences and introducing autoregressive transformer architectures tailored to this sequential representation. Specifically, MeshGPT [123] leverages VQVAE [124] to learn a mesh vocabulary and employs a decoder-only transformer to autoregressively generate triangle meshes. To mitigate cumulative errors inherent in VQVAE-based methods, MeshXL [125] further introduces a neural coordinate field for sequential 3D mesh representation. To address the problem that the above methods struggle to learn the shape and topology distributions together, MeshAnything [126] uses a pre-trained encoder [127] to inject shape features into a VQVAE-based sequence, removing the need to learn shape distribution and enabling focused topology learning.
@@ 2.4.2 Occupancy
Occupancy [128, 42] refers to the property that describes whether a given point in a 3D space is inside or outside a surface or object. Several methods have been proposed to achieve feed-forward occupancy prediction from images. For example, Any-Shot GIN [129] aims to model occupancy-based 3D implicit reconstruction. It begins with front-back depth estimation to generate depth maps for constructing a voxel-based representation, subsequently extracting 3D features from this volume to infer the occupancy of any 3D point in the space; MCC [130] encodes a compressed representation of the scene appearance and geometry, and then uses the representation to predict occupancy probability and RGB color for each 3D point; ZeroShape [131] employs intermediate geometric representation and explicit reasoning to achieve 3D occupancy regression. While there are also many works on semantic occupancy prediction for autonomous driving, we do not cover these methods here, as the focus of this survey is on 3D reconstruction and view synthesis.
@@ 2.4.3 SDF
Signed Distance Function (SDF) [132] is a mathematical function that represents the geometry of a shape or surface in space. For any point in 3D space (or 2D), the SDF returns the shortest distance from that point to the surface of the object. The sign of the distance indicates whether the point is inside or outside the object. Several methods have been proposed to enable feed-forward SDF representations. For example, SparseNeuS [118] initially builds a hierarchy of volumes that represent local surface details, which are then used to infer SDF-based surfaces through a progressive coarse-to-fine process; Shap-E [25] transforms point clouds and RGBA input images into a sequence of latent vectors for SDF prediction. Some works also incorporate multiview stereo (MVS) for SDF prediction. For instance, C2F2NeUS [133] constructs a hierarchy of geometric frustums to capture local-to-global geometry for SDF prediction; UFORecon [134] introduces a cross-view matching transformer to extract cross-view matching features to construct hierarchical correlation volumes, enabling SDF-based reconstruction with limited camera view overlaps. Without relying on MVS, CRM [135] predicts SDF by incorporating geometric priors into network designs based on the spatial alignment between triplanes and six input orthographic views generated by a multiview diffusion model. There have also been works that explore transformers for SDF prediction, including VolRecon [136] which employs a view transformer to integrate cross-view features and a ray transformer to estimate SDF values, and ReTR [137] which introduces an occlusion transformer and a render transformer to fuse features and perform rendering.
Figure 6: Typical frameworks of regression-based 3D-free models. The image is adapted from  [7].
@@ 2.5.1 Regression-based View Synthesis
Scene representation transformer (SRT) [138] leverages a transformer-based encoder to map multiview input images to latent representations first and then outputs novel-view images from a transformer-based decoder with light field rays. RUST [139] inherits an encoder-decoder architecture and enables novel view synthesis solely from RGB images, without the need for camera poses. With a focus on object-centric 3D scenes, OSRT [140] incorporates a slot attention module on SRT to map the encoded latent representations to object-centric slot representations. To extend SRT to large-scale scenes, RePAST [141] integrates relative camera pose information into the attention layer of SRT. However, these methods often suffer from degraded details and suboptimal rendering quality. To address this issue, several approaches incorporate geometric information to improve model performance. For example, GPNR [142] integrates epipolar geometry within its encoder-decoder architecture, while Du et al. [143] introduce a multiview vision transformer and epipolar line sampling to improve scene geometry. GBT [144] incorporates ray distance-based geometry reasoning into multihead attention layers of transformers in the encoder and decoder. GTA [145] introduces geometric transform attention to embed the geometrical structure of tokens into the transformer and integrates it into SRT to enhance transformer-based rendering. However, despite the improved model performance, geometrical designs often integrate additional 3D inductive biases. LVSM [7] removes the geometrical designs and leverages a transformer-based large reconstruction model with self-attention to regress the target view pixels. To enable novel view synthesis without ground truth of cameras, RayZer [146] introduces a self-supervised multiview model that first learns camera parameters and latent scene representations from unposed input images, and then renders novel views.
Figure 7: Representative frameworks of generative 3D-free models. The samples are adapted from  [147] and  [148].

### v4

@@ 2.3.1 Gaussian Map
A Gaussian map refers to a 2D-based representation of 3D Gaussians, where each 2D location encodes a 3D Gaussian. As a pioneering effort, Splatter Image [SRV24] employs a U-Net encoder-decoder architecture [RFB15] to predict pixel-aligned 3D Gaussians for single-view 3D object reconstruction as illustrated in Fig. 4. To improve the reconstruction quality, subsequent methods leverage generic scene priors learned from large-scale datasets. GRM [XSY24] directly maps input image pixels to pixel-aligned 3D Gaussians for feed-forward object reconstruction, based on LRM [HZG23] that learns general reconstruction priors from large-scale datasets of 3D objects; Flash3D [SIZ24] further introduces a high-quality depth predictor as prior to achieve single-view scene-level reconstruction. Concurrently, GS-LRM [ZBT24] incorporates Transformer-based LRM to formulate per-pixel Gaussian prediction as a sequence-to-sequence mapping, achieving remarkable performance across both objects and scenes. This line of research has been further advanced by several works: eFreeSplat [MLSY24] leverages a large vision transformer encoder [DBK20] as 3D priors for Gaussian image prediction; Long-LRM [ZTZ24] that combines Mamba2 blocks [DG24] with Transformer layers to handle long sequences of input images; and FreeSplatter [XGS24] achieves pose-free Gaussian map prediction. However, these works are limited to reconstructing existing image observations without generative capabilities. To overcome this limitation, LGM [TCC24] introduces pre-trained diffusion models [LGL24, SCZ23, SWY23, WS23] to generate multiview images, which are then used by large multiview Gaussian models for multiview Gaussian prediction. Wonderland [LCG25] further enhances this approach by leveraging a pre-trained video diffusion model [YTZ24] to generate informative video latents from a single image for 3DGS prediction. In addition to the methods above, another line of work focuses on the geometric quality of 3D scene reconstruction by incorporating geometric designs or priors, such as epipolar constraints, cost volumes, and pre-trained reconstruction models.
@@
Epipolar-based Methods. As a pioneering epipolar-based method, PixelSplat [CLTS24] leverages an epipolar line to resolve the scale ambiguity issue and capture cross-view features, estimating a probabilistic depth distribution as 3D Gaussian positions. Although PixelSplat is effective in regions strongly correlated with the input observations, it struggles in areas of high uncertainty, leading to blurry or failed reconstructions. To address this, LatentSplat [WRI24] combines the strengths of regression-based and generative approaches to obtain high-quality reconstructions in uncertain areas. Building further on PixelSplat, GGRt [LGZ24] introduces a joint learning framework that integrates pose estimation with a 3DGS prediction network to enable pose-free 3D Gaussian prediction.
@@
Cost Volume-based Methods. A key limitation of PixelSplat is the inherent ambiguity and unreliability in mapping image features to depth distributions, resulting in suboptimal geometry reconstruction. To address this problem, MVSplat [CXZ24] and MVSGaussian [PMSB24] adopt a plane-sweeping-based cost volume to facilitate multiview Gaussian image prediction, leveraging cross-view feature similarities to improve depth estimation. However, these methods heavily depend on precise multiview feature matching, which becomes particularly challenging in scenes with occlusions, low texture, or repetitive patterns. To address this issue, TranSplat [ZZL25] introduces a depth-aware deformable matching transformer to generate a depth confidence map to facilitate multiview feature matching. Similarly, DepthSplat [XPW24] uses a multiview depth model that leverages pre-trained monocular depth features to enhance the feed-forward 3DGS reconstruction. This line of research is further extended by several works. HiSplat [TYY24] introduces hierarchical representations to capture large-scale structures and fine texture details; PanSplat [ZXW24] builds a hierarchical spherical cost volume for 4K panorama view synthesis; MVSplat360 [CZX24] adapts MVSplat to support 360° novel view synthesis for large-scale scenes; and LongSplat [HWG25] designs an online updating framework for long-sequence image streams.
@@
Pre-trained Reconstruction Models. With the advent of feed-forward pointmap reconstruction methods [WLC24a, LCR24, WA24] as mentioned in Sec. 2.2, several feed-forward 3DGS methods directly use pre-trained models to generate dense pointmaps for 3D Gaussian reconstruction. For example, Splatt3R [SZLP24] builds on the large-scale pretrained foundation 3D MASt3R model [LCR24] by integrating a Gaussian decoder, enabling pose-free feed-forward 3DGS. NoPoSplat [YLX24] also uses MASt3R as the backbone and predicts 3D Gaussians in a canonical space without ground-truth camera poses and depth. SmileSplat [LFTL24] instead uses DUSt3R as the backbone to predict Gaussian surfels with a multi-head Gaussian regression decoder. SelfSplat [KYP24] unifies Gaussian prediction with self-supervised learning of depth and camera poses, enabling simultaneous prediction of geometry, pose, and Gaussian attributes. However, methods relying on DUSt3R and MASt3R inherit the limitation of pairwise inputs, which restricts their scalability. Thus, PREF3R [CYY24] adopts Spann3R [WA24] as the geometric prior, which introduces a spatial memory network to handle multiview images. Except for feed-forward pointmap prediction models, pre-trained 3D diffusion models [JN23] have also been explored as a geometric prior for 3D Gaussian prediction [LGD24].
@@ 2.3.2 Gaussian Volume
Gaussian volume [CXE24] represents 3D with Gaussian voxel grids, where each voxel comprises multiple Gaussian primitives. A typical feed-forward 3DGS method using a Gaussian volume representation is LaRa [CXE24], which aims to reduce the heavy training cost associated with 360° bounded radiance field reconstruction. As shown in Fig. 4 (b), it first builds 3D features and embedding volumes and then leverages a volume transformer to reconstruct a Gaussian volume, enabling progressively and implicitly feature matching that leads to higher quality results and faster convergence. Building on the notion of structured volumetric Gaussians, GaussianCube [ZCY24] proposes a structured and explicit radiance representation for 3D object generation from a single image. In parallel, QuickSplat [LHND25], based on G3R [CWY24], predicts voxel-level features using neural networks, which are subsequently decoded into Gaussian primitives with MLPs for surface reconstruction. To further enrich volumetric representations, GD [NSK25] builds upon LaRa and introduces a generative densification that exploits prior knowledge from large multiview datasets and densifies feature representations from feed-forward 3DGS. SCube [RLL24] advances this direction toward large-scale scene reconstruction by proposing VoxSplat, a high-resolution sparse-voxel Gaussian representation generated via a hierarchical latent diffusion model conditioned on sparse posed images. Most recently, AnySplat [JMX25] constructs volumetric Gaussians by voxelizing the Gaussian maps from variable numbers of input views, yielding a unified and efficient model capable of handling both sparse and dense views.
To mitigate the high computational cost of the 3D volume, an efficient triplane structure [PNM20, CLC22] is also explored for Gaussian volume prediction. It is typically constructed by predicting a triplane representation first and then leveraging the latent triplane features to decode 3D Gaussians, as illustrated in Fig. 4. For example, Triplane-Gaussian [ZYG24] leverages several transformer-based networks pre-trained in large-scale datasets to build a Gaussian triplane, enabling high-quality single-view 3D reconstruction. AGG [XYM24] also mixes triplane and 3D Gaussians, which first represents scene textures as triplane and then decodes 3D Gaussians from triplane-based texture features queried by 3D locations.
@@ 2.4.1 Mesh
Meshes are compatible with various graphics pipelines and have gained significant attention in feed-forward 3D reconstruction in recent years. For instance, Pixel2Mesh [WZL18] generates a 3D mesh from a single image by extracting features with a 2D CNN and progressively deforming an initial mesh; Mesh R-CNN [GMJ19] introduces a voxel branch based on Mask R-CNN [HGDG17] to predict a coarse cubified mesh followed by further refinement.
This field has been further advanced by the development of diffusion models. For example, One-2-3-45 [LXJ23] builds on the diffusion-based model to generate multiview images, which are then processed by a generalizable surface reconstruction module [LLW22] for mesh reconstruction; One-2-3-45++ [LSC24] further utilizes a 3D diffusion-based module conditioned on multiple views to generate a textured mesh in a coarse-to-fine manner. To improve geometry consistency, Wonder3D [LGL24] introduces a cross-domain diffusion model to generate multiview-consistent normal maps and RGB images simultaneously, which can be used for mesh reconstruction. Similarly, Unique3D [WLC24b] employs a multiview diffusion model alongside a normal diffusion model to generate multiview-consistent images and normal maps, which are fed into a fast and consistent mesh reconstruction module.
In addition, several methods are proposed to take advantage of large reconstruction model [HZG23] to achieve high-quality mesh reconstruction. For example, MeshLRM [WZB24] integrates differentiable surface extraction and rendering into a large reconstruction model, enabling direct 3D mesh generation; InstantMesh [XCG24] utilizes a transformer-based large reconstruction model to generate a high-quality 3D mesh from multiview images; MeshFormer [LZW24a] leverages 3D voxel representations and combines transformers with 3D convolutions to improve 3D mesh geometry.
To generate artist-created meshes with high-quality topology, several methods draw inspiration from large language models, treating 3D meshes as sequences and introducing autoregressive transformer architectures tailored to this sequential representation. Specifically, MeshGPT [SAA24] leverages VQVAE [VDOV17] to learn a mesh vocabulary and employs a decoder-only transformer to autoregressively generate triangle meshes. To mitigate cumulative errors inherent in VQVAE-based methods, MeshXL [CCP24] further introduces a neural coordinate field for sequential 3D mesh representation. To address the problem that the above methods struggle to learn the shape and topology distributions together, MeshAnything [CHH24] uses a pre-trained encoder [ZLC23] to inject shape features into a VQVAE-based sequence, removing the need to learn shape distribution and enabling focused topology learning.
@@ 2.4.2 Occupancy
Occupancy [MON19, PNM20] refers to the property that describes whether a given point in a 3D space is inside or outside a surface or object. Several methods have been proposed to achieve feed-forward occupancy prediction from images. For example, Any-Shot GIN [XCB22] aims to model occupancy-based 3D implicit reconstruction. It begins with front-back depth estimation to generate depth maps for constructing a voxel-based representation, subsequently extracting 3D features from this volume to infer the occupancy of any 3D point in the space; MCC [WJM23] encodes a compressed representation of the scene appearance and geometry, and then uses the representation to predict occupancy probability and RGB color for each 3D point; ZeroShape [HST24] employs intermediate geometric representation and explicit reasoning to achieve 3D occupancy regression. While there are also many works on semantic occupancy prediction for autonomous driving, we do not cover these methods here, as the focus of this survey is on 3D reconstruction and view synthesis.
@@ 2.4.3 SDF
Signed Distance Function (SDF) [PFS19] is a mathematical function that represents the geometry of a shape or surface in space. For any point in 3D space (or 2D), the SDF returns the shortest distance from that point to the surface of the object. The sign of the distance indicates whether the point is inside or outside the object. Several methods have been proposed to enable feed-forward SDF representations. For example, SparseNeuS [LLW22] initially builds a hierarchy of volumes that represent local surface details, which are then used to infer SDF-based surfaces through a progressive coarse-to-fine process; Shap-E [JN23] transforms point clouds and RGBA input images into a sequence of latent vectors for SDF prediction. Some works also incorporate multiview stereo (MVS) for SDF prediction. For instance, C2F2NeUS [XGW23] constructs a hierarchy of geometric frustums to capture local-to-global geometry for SDF prediction; UFORecon [NKH24] introduces a cross-view matching transformer to extract cross-view matching features to construct hierarchical correlation volumes, enabling SDF-based reconstruction with limited camera view overlaps. Without relying on MVS, CRM [WWC24] predicts SDF by incorporating geometric priors into network designs based on the spatial alignment between triplanes and six input orthographic views generated by a multiview diffusion model. There have also been works that explore transformers for SDF prediction, including VolRecon [RWZ23] which employs a view transformer to integrate cross-view features and a ray transformer to estimate SDF values, and ReTR [LHC23] which introduces an occlusion transformer and a render transformer to fuse features and perform rendering.
Figure 5: Typical frameworks of regression-based 3D-free models. The image is adapted from  [JJT25].
@@ 2.5.1 Regression-based View Synthesis
Scene representation transformer (SRT) [SMP22] leverages a transformer-based encoder to map multiview input images to latent representations first and then outputs novel-view images from a transformer-based decoder with light field rays. RUST [SMK23] inherits an encoder-decoder architecture and enables novel view synthesis solely from RGB images, without the need for camera poses. With a focus on object-centric 3D scenes, OSRT [SDM22] incorporates a slot attention module on SRT to map the encoded latent representations to object-centric slot representations. To extend SRT to large-scale scenes, RePAST [SDS23] integrates relative camera pose information into the attention layer of SRT. However, these methods often suffer from degraded details and suboptimal rendering quality. To address this issue, several approaches incorporate geometric information to improve model performance. For example, GPNR [SESM22] integrates epipolar geometry within its encoder-decoder architecture, while Du et al. [DSTS23] introduce a multiview vision transformer and epipolar line sampling to improve scene geometry. GBT [VAST23] incorporates ray distance-based geometry reasoning into multihead attention layers of transformers in the encoder and decoder. GTA [MJWG23] introduces geometric transform attention to embed the geometrical structure of tokens into the transformer and integrates it into SRT to enhance transformer-based rendering. However, despite the improved model performance, geometrical designs often integrate additional 3D inductive biases. LVSM [JJT25] removes the geometrical designs and leverages a transformer-based large reconstruction model with self-attention to regress the target view pixels. To enable novel view synthesis without ground truth of cameras, RayZer [JTW25] introduces a self-supervised multiview model that first learns camera parameters and latent scene representations from unposed input images, and then renders novel views.
Figure 6: Representative frameworks of generative 3D-free models. The samples are adapted from  [LWVH23] and  [LSW24].